OpenStudy (anonymous):

Pick a two-digit number greater than 25. Rewrite your two-digit number as a difference of two numbers. Show how to use the identity (x − y)2 = x2 − 2xy + y2 to square your number without using a calculator. I picked 26 30-4=26 division 26^2=(30-4)^2= 30^4*2*30-4+4^2 My teacher told me I need to FOIL. Which I don't understand what to do.. IF you could show me how with steps that would be great.

OpenStudy (anonymous):

The letters FOIL stand for First, Outer, Inner, Last. First means multiply the terms which occur first in each binomial. Then Outer means multiply the outermost terms in the product. Inner means multiply the innermost two terms. Last means multiply the terms which occur last in each binomial. Then simplify the products and combine any like terms which may occur.

OpenStudy (anonymous):

@drumminboy918 So could you help me with that, Which equation do I use FOIL on?

OpenStudy (anonymous):

26^2=(30-4)^2= 30^4*2*30-4+4^2 This is the Equation you use it on

OpenStudy (anonymous):

@drumminboy918 What about 30^4*2*30-4+4^2 the -4+4^2, do I not worry about it and just do 30^4*2*30 using foil

OpenStudy (anonymous):

You use it on both sides of the equation I told you to do

OpenStudy (anonymous):

@drumminboy918 I don't understand, how do I multiply 26^2=(30-4)^2, can you help me

OpenStudy (anonymous):

Here I'll give you a example

OpenStudy (anonymous):

Example: (x + 2)(x + 5) = x·x + x·5 + 2·x + 2·5 First Outer Inner Last = x2 + 7x + 10

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do the same thing here 26^2=(30-4)^2= 30^4*2*30-4+4^2

OpenStudy (anonymous):

What did you multiply to get x2+7x+10 in the example

OpenStudy (anonymous):

I am just simplifying the left side

OpenStudy (anonymous):

Can you give me the part of my equation to simplify for the first part of FOIL

OpenStudy (anonymous):

I man right

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

mean* gosh I cant type today

OpenStudy (anonymous):

so 4^2 is my right side do I simply that

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So 16 is the answer for F, then we move on to O?

OpenStudy (anonymous):

First, Outer, Inner, Last

OpenStudy (anonymous):

Your getting the hang of it

OpenStudy (anonymous):

Wouldn't the outermost be the right side to?

OpenStudy (anonymous):

The outer of the right side

OpenStudy (anonymous):

I just did that?

OpenStudy (anonymous):

all your doing is simplifying

OpenStudy (anonymous):

that's all FOIL is

OpenStudy (anonymous):

A fancy way of simplifying

OpenStudy (anonymous):

So can you tell me what part I neeed to simplify for O now? as 4^2 is the outtermost

OpenStudy (anonymous):

Sadly I cant do this for you, but what ive seen so far, you can do this. But this place is easy simple and fun http://www.algebrahelp.com/lessons/simplifying/foilmethod/

OpenStudy (anonymous):

@drumminboy918 but is what I got for the first part right? I'm not asking for the answers just what you are telling me is going against what you told me above.

OpenStudy (anonymous):

Im not sure, that why. I use this link all the time

OpenStudy (anonymous):

@drumminboy918 Yeh, i'm looking at the site, so 16 is not correct.. I don't understand how I use foil on this if I don't have 2 sets of parentheses

OpenStudy (anonymous):

You need 2 sets of parentheses. Choose a different equation with a set of 2 parentheses.

OpenStudy (anonymous):

@Klip My teacher told me the equation was right for this and that I just needed to foil, so what I did was I did this 26^2=(30-4)^2= 30^4*2*30-(4+4)^2. Then I foiled

OpenStudy (anonymous):

@Klip So that would work? Thanks